# Whether to trust: the ML leap of faith

### v1

@@ 3 Proposed architecture
Model-architecture functional compartmentalisation enables user oversight over data and objective function, increasing human agency. We use an expert-validated rules-based AI model to create a verified point of reference that can be tested a priori by ML model users. Our LoF matrix visually contrasts this reference standard to our ML model’s output, using the same data and objective function. Areas of agreement can be readily trusted and what is unique to the ML model can be identified and addressed. We propose three simple objective metrics to measure demonstrated trust, distinguishing intention-setting and follow-through, and whether trust was deserved.
@@ 3.1 Enabling human agency and oversight
Model-architecture functional compartmentalisation is used to separate data and objective function from model reasoning. Each user was responsible for pre-validating their own data. This exercise uses text and data visualisation techniques, extending the text-driven interpretable-feature-space approach of Zytek et al. [19] to include target variables and the addition of a visual dimension, to make the data more accessible to users who benefit cognitively from visual processing of data.
Validation started with data quality – percentage of each day with complete data for each device – which the system reported daily and weekly against a pre-briefed format and target (70%). Participants progressed to baseline, the next phase of the study, if their one-week trial data was over target.
Participants then assessed whether the data collected during their trial week reflected their perception. Each feature input was provided in a text and visual format that had been reviewed by sleep experts and by 35 participants to ensure each was readable, human-worded and understandable, e.g.: ‘You consume 152mg of caffeine a day (0-286mg range)’, and the chart on the left-hand side of Figure 3.
Figure 3: Caffeine data validation example: upfront data-quality trial chart on the left; baseline chart on the right. The rules embedded in the rules-based model were used to colour-code the latter – in this example, blue for <230mg; red for >400mg; yellow for in-between. Blue was used, rather than green, to make the charts accessible to colour blind participants.
Participants were educated on how to read each chart, what the units meant, and how they were calculated. Participants were encouraged to challenge data accuracy. Some data required considerable interrogation and education because the data and/ or the units were unfamiliar, especially if they were asleep when the data was recorded (e.g. temperature and illumination, as shown in Figure 4).
@@
Participants trialed different behaviours for unfamiliar items to see how the charts responded. Some changed how an input was recorded if the issue was a diary-input mistake. These follow-ups continued until participants were comfortable their data was accurate.
Participants then reviewed at least 1 month of complete, quality days of baseline data, to confirm target variables and feature inputs were accurate. This included text and charts with sleep-expert-validated logic from the rules-based model, e.g.: ‘You had 330mg of caffeine/ day (blue line). You were over 230mg, the amount which impacts the average person’s sleep, on 82% of days (yellow); over NHS’s 400mg recommended max on 11% of days (red).’ and the chart on the right-hand side of Figure 3.
Each participant had unique sleeping challenges (e.g. waking in the night or too early or having trouble sleeping at night) and objectives: some already achieved well over 7 hours’ sleep and just wanted to improve SWS; some only got 6 hours of sleep and wanted to focus on TST. Sleep objectives were set after the 1-week trial and confirmed once they validated their baseline data and were educated on what was realistic. They thus set model objective function, increasing their agency.
Participants were given the opportunity to remove any baseline data that they did not believe was accurate, but none did. This is an example of demonstrating trust in the data component of an ML model, and of a non-expert being able to build intrinsic trust in part of a complex ML model, because they are able to understand the data inputs and objective function, albeit with support. Data validation, with custom question responses, demonstrated personalisation, an immediacy behaviour, and increased model trustworthiness by ensuring correct inputs.
@@ 3.2 Discerning the ML leap of faith
The system recommendation was positioned and visually depicted as an intervention ‘menu’ using a LoF matrix which, as shown in Figure 5, discerns the leap of faith required to trust the ML algorithms. By using a separate a priori expert-validated rules-based AI model to create a reference standard, we contrast the ML models’ prioritisation of which sleep-hygiene changes would make the most difference to the user’s sleep objective, with what an expert would conclude from the same inputs. The user is able to see where the models agree, and where they do not.
Distance from best practice, assessed by the rules-based model, is used to prioritise opportunities on the vertical axis; ML-model feature importance is used to prioritise them on the horizontal. The further away an individual was on best practice, the more impact changing that factor could have on the individual’s sleep. The rules-based model assigned 1 of 3 levels of opportunity to each intervention, or a 4th category: ‘angel’, meaning they had achieved best practice. The 35 design-phase participants preferred 3 levels of opportunity, associating them with a ‘good, better, best’ paradigm, and appreciated the self-efficacy boost of knowing they were already achieving best practice.
Four demarcations were also used to categorise ML opportunities using feature importance, measured through average gain across all splits where a feature was used. The higher a feature’s importance, the higher an individual’s sensitivity to that feature. Feature-importance thresholds were set so each individual had at least 4 opportunities in the last 2 columns of the LoF matrix, equivalent to 2- and 3-star rules-based opportunities. The ML model was more discriminating, e.g. double the number of 3-star opportunities in Figure 5, and 9 ‘angels’, compared to 1 ‘angel’ in the rules-based model.
Participants were informed that the vertical axis was the expert conclusion based on their data and objective, and the horizontal axis was the ML assessment of their individual sensitivity using the same inputs. The supporting rationale, combining the individual’s validated data with sleep-research insights, was displayed in the user interface, using the rules-based model logic.
Participants were instructed to select interventions that they: understood; were able to control; and were motivated to change [43]. They were required to give themselves an ‘adequate sleep opportunity’ for 7 hours of sleep, based on SE (of time asleep while in bed). They were given decisional freedom to choose up to 3 other interventions on the LoF matrix. Once they chose their interventions, they were asked to set specific targets (e.g. get up between 6:45 and 7:15 every day, including weekends).
Figure 6: Example tracker daily text and dashboard.
The user-interface tracker provided daily feedback over the course of 1 month of intervention on how participants were doing against these actions. The user-interface tracker presented a simple dashboard accompanied by an automated text message each day. The user could investigate each metric, using the same data visualisation as they had used to validate their data (now colour-coded relative to their target) and longer-term views of the changes.
@@ 3.3 Measuring the ML leap of faith
Intervention choice, compliance, and sleep outcomes were recorded, enabling three relative-trust metrics to be assessed: DIRTI, DAFTI, and DOTI. We illustrate each using pilot data; see Figure 7 for how these metrics relate to the types of trust discussed in the literature.
@@
Figure 11: DAFTI: average ML priority after follow-through (dark blue) divided by average rule priority (light blue) yields the relative trust demonstrated in ML after follow-through (mid-blue). If the participants showed any improvement vs baseline, the score was included. Blue lines are averages.
Our ML models were designed to be trustworthy, so the outstanding deserved-trust question is whether there is a relationship between the models’ prioritisation of recommendations and the outcomes delivered, i.e. did participants who implemented high-scoring interventions get better outcomes? Was the risk worth it? The Deserving Of Trust index (DOTI)measures whether the trust was worth it, reflecting the constraints faced once the participant acted. There will be a placebo effect at play, which can be considerable, but there is no reason to assume it would not align to demonstrated trust.
Demonstrated trust is shown on an absolute basis on the left of Figure 12 – i.e. the correlation (${r}^{2}$) between each individual’s average ML or rules score and their percentage increase in TST and SWS. We compare baseline to the last seven days of intervention because most participants worked towards target behaviour in stages over four weeks of intervention. In this pilot, the strongest relationship was for ML-interventions at follow-through (dark blue): TST ${r}^{2}$ increases from 0.58 at intention to 0.76; SWS from 0.22 to 0.54. An increase is consistent with a trustworthy model because recommendations only work if people act on them. Relative DOTI, the ratio of the absolute ML score and the rules score, is shown on the right. Trust was more deserved in the ML algorithm for TST than for SWS, which may have been influenced by lower SWS model accuracy. At follow-through, our ML model deserved more trust than the rules-based one; had only intention been considered, the SWS rules-based model would have (incorrectly) been seen to deserve more trust.
Figure 12: Absolute DOTI on the left: TST at top and SWS on bottom: correlation (${r}^{2}$) between intention and follow-through scores, and % improvement in last 7 days’ sleep end points relative to baseline. Relative DOTI is on the right, calculated by dividing the ML metric by the rules-based one.
@@ 4 Discussion
The proposed architecture allows the ML leap of faith to be discerned and measured, and therefore to be managed over time. Experimental design lessons include:
More human agency and control is afforded by data validation and objective-function specification, enabled by separating data and objective function from model reasoning and data visualisation. Users build intrinsic trust by pre-validating data in an accessible format, and deciding the model objective.
@@
An explicit reference standard, created through a collectively-developed rules-based model, enables experts to test it against their mental model a priori. They should do so until they intrinsically trust it, so it becomes a verified point of reference. Industry or regulatory bodies could maintain the standard as critical infrastructure – e.g. for assurance audits or to train up new experts. Maintenance should include regularly stress testing it to keep assumptions and use cases up-to-date, and verify ongoing trust. Many domains already use rules-based models, which should evolve into a reference standard.
@@
A LoF matrix visually and accessibly contrasts the output of the reference model with the ML model’s a posteriori output, using the same data and objective function, thus bringing the ML leap of faith into sharp focus. This provides real transparency to model users: they should be inclined to trust where models agree, and understand a leap of faith is only required where they disagree. Experts can stress test the areas of disagreement, updating the rules-based model with new insights or to capture drift, thus narrowing the leap of faith as an aspect of operations management. Different ML algorithms will require different leaps of faith, because a priori levels of expert insight differ, as do model accuracies.
@@
By being explicit about the ML leap of faith, we can assess demonstrated and deserved trust. For the first time, our metrics measure trust in terms of actions and outcomes. If trust is deserved, so relative DOTI increases over time; DIRTI and DAFTI should also increase. These simple metrics allow trust comparisons across models, and flag the need for investigation if they trend down. Industry regulators could track companies’ ML-model-trust metrics, especially DOTI, to understand trends. Continuous audit processes using these metrics would preserve ML algorithm IP whilst building public assurance.
Limitations
@@
The matrix and metrics may be perceived as insufficiently technical. They are accessible by design, so they can be used and understood by experts, businesses, and regulators as well as computer/ data scientists. They need to be applied at scale to reach empirical conclusions about trust for different users/ domains/ situations, and what drives it, objectively. The proposed metrics do not measure attitudinal trust or model trustworthiness. Instead they measure how attitude is reflected in decisions, actions, and outcomes, which drive ML adoption. Because our proposed metrics only measure intentions, actions and outcomes, they could be applied to an untrustworthy system and used to encourage adoption. If GDPR (or similar) is not followed, trust metrics could be inappropriately sold and/ or used to discriminate against people based on their general or specific trust inclination.
@@
Participant time for data validation can be substantial if data does not readily align with perception. Our approach may be most suitable for professionals who use model recommendations often, so their time commitment pays off as they focus on where there is a leap of faith to be overcome. The time required to develop a reference standard from scratch can be considerable, requiring: substantial programmer subject-matter education; individuals who can translate expert knowledge into rules; and time from experts on each iteration, and to test the resulting model for different use cases. However, experts may resist allocating this time to ML, especially if they have a high opinion of their own capability and/or are concerned about their roles being replaced or diminished. The data-validation approach is not universally accessible: our participants had at least one university degree and understood complex, novel concepts and images. Work is needed to evaluate accessibility for other levels of educational attainment. Prioritisation, unlike direct regression/ classification comparisons, requires calibration if it is to be included in a LoF matrix: different approaches should be tested.
@@ Conclusion
The proposed architecture is a more practical, rigorous way of building trust in an ML model than trying to explain it to people who have a different (probably unarticulated) mental model. Critical novel architecture features include: user data validation and objective-function specification; an explicit reference standard; a LoF matrix; and demonstrated- and deserved-trust metrics.

### v2

@@ 4 Proposed architecture
Human systems of trust, in high-stakes environments like medicine, rely on trusted colleagues pressure-testing conclusions based on validated inputs and standard practices, given desired patient outcomes [43]. Our participants validated input data and set the agent’s objectives; experts tested whether our rules-based AI model reflected their mental model: they interrogated the logic with our validated, to test whether their standard practices had been correctly implemented. This model-architecture functional compartmentalisation therefore enables user oversight over data and objective function, increasing their agency. Our expert-validated rules-based AI model creates a verified reference agent that can be tested a priori by ML model users. Our LoF matrix visually contrasts this reference agent to our ML agent’s recommendation, using the same data and objective function. Areas of agreement can be readily trusted and what was unique to the ML agent can be identified and addressed. We propose three simple objective metrics to measure demonstrated trust, distinguishing intention-setting and follow-through, and whether trust was deserved.
@@ 4.1 Enabling human agency and oversight
Model-architecture functional compartmentalisation is used to separate data and objective function from model reasoning. Each user was responsible for pre-validating their own data. This exercise uses text and data visualisation techniques, extending the text-driven interpretable-feature-space approach of Zytek et al. [15] to include target variables and adds data visualisation, to make the data more accessible to users who benefit cognitively from visual processing.
Validation started with data quality – percentage of each day with complete data for each device – which a data-quality agent reported daily and weekly against a pre-briefed target (70%). Participants progressed to baseline, the next phase of the study, if their one-week trial data was over target. Participants then assessed whether the data collected during their trial week reflected their perception. Each feature input was provided in a text and visual format that had been reviewed by sleep experts and by 35 participants to ensure each was readable, human-worded and understandable, e.g.: ‘You consume 152mg of caffeine a day (0-286mg range)’, and the chart on the left-hand side of Figure 3.
Figure 3: Caffeine data validation example: upfront data-quality trial chart on the left; baseline chart on the right. The rules embedded in the rules-based model were used to colour-code the latter – in this example, blue for <230mg; red for >400mg; yellow for in-between.
@@
Participants were educated on how to read each chart, what the units meant, and how they were calculated. Participants were encouraged to challenge data accuracy. Some data required considerable interrogation and education because the data and/ or the units were unfamiliar, especially if they were asleep when the data was recorded (e.g. temperature and illumination, as shown in Figure 4). Participants trialed different behaviours for unfamiliar items to see how the charts responded. Some changed how an input was recorded if the issue was a diary-input mistake. These follow-ups continued until participants were comfortable their data was accurate. Participants then reviewed at least 1 month of complete, quality days of baseline data, to validate that the target variables and feature inputs were accurate. This included text and charts with sleep-expert-validated logic from the rules-based agent, e.g.: ‘You had 330mg of caffeine/ day (blue line). You were over 230mg, the amount which impacts the average person’s sleep, on 82% of days (yellow); over NHS’s 400mg recommended max on 11% of days (red).’ and the chart on the right-hand side of Figure 3.
Each participant had unique sleeping challenges (e.g. waking in the night or too early or having trouble sleeping at night) and objectives: some already achieved well over 7 hours’ sleep and just wanted to improve SWS; some only got 6 hours of sleep and wanted to focus on TST. Sleep objectives were set after the 1-week trial and confirmed once users validated their baseline data and were educated on what was realistic. They thus set the agent’s objective function, increasing their agency. Participants were given the opportunity to remove any baseline data that they did not believe was accurate, but none did. This is an example of demonstrating trust in the data component of an ML model, and of a non-expert being able to build intrinsic trust in part of a complex ML model, because they are able to understand the data inputs and objective function, albeit with support. Data validation and custom question responses, is an example of personalisation, an immediacy behaviour, also increasing agent trustworthiness by ensuring correct inputs.
@@ 4.2 Discerning the ML leap of faith
The recommendation was visually depicted as an intervention ‘menu’ using a LoF matrix which, as shown in Figure 5, discerns the leap of faith required to trust the ML agents. By using a separate a priori expert-validated rules-based AI agent, we contrast the ML agents’ prioritisation of which SH changes would make the most difference to the user’s sleep objective, with what an expert would conclude from the same inputs. The user is able to see where the agents agree, and where they do not. Distance from best practice, assessed by the rules-based agent, is used to prioritise opportunities on the vertical axis; ML-agent feature importance is used to prioritise them on the horizontal.
@@
The further away an individual was on best practice, the more impact changing that factor could have on the individual’s sleep. The rules-based agent assigned 1 of 3 levels of opportunity to each intervention, or a 4th category: ‘angel’, meaning they had achieved best practice. The 35 design-phase participants preferred 3 levels of opportunity, associating them with a ‘good, better, best’ paradigm, and appreciated the self-efficacy boost of knowing they were already achieving best practice. Four demarcations were also used to categorise ML opportunities using feature importance, measured through average gain across all splits where a feature was used. The higher a feature’s importance, the higher an individual’s sensitivity to that feature. Feature-importance thresholds were set so each individual had at least 4 opportunities in the last 2 columns of the LoF matrix, equivalent to 2- and 3-star rules-based opportunities. The ML agent was more discriminating, e.g. double the number of 3-star opportunities in Figure 5, and 9 ‘angels’, compared to 1 ‘angel’ in the rules-based model. Participants were informed that the vertical axis was the expert conclusion based on their data and objective, and the horizontal axis was the ML assessment of their individual sensitivity using the same inputs. The supporting rationale, combining the individual’s validated data with sleep-research insights, was displayed in the user interface, using the rules-based model logic.
Participants were instructed to select interventions they: understood; were able to control and motivated to change [44]. They were required to give themselves an ‘adequate opportunity’ for 7 hours of sleep, based on SE (of time asleep while in bed). They were given decisional freedom to choose up to 3 other interventions on the LoF matrix. Once they chose their interventions, they set specific targets (e.g. get up between 6:45 and 7:15 every day, including weekends). The user-interface tracker provided daily feedback over the course of 1 month of intervention on how participants were doing against these actions. The user-interface tracker was a simple dashboard accompanied by an automated text message each day. The user could investigate each metric, using the same data visualisation as they had used to validate their data (now colour-coded relative to their target).
@@ 4.3 Measuring the ML leap of faith
Intervention choice, compliance, and sleep outcomes were recorded, enabling three relative-trust metrics to be assessed: DIRTI, DAFTI, and DOTI. We illustrate each using pilot data. See Figure 6 for how these metrics relate to the types of trust discussed in the literature, and Box 4.3 for formulae.
@@
Figure 10: DAFTI: average ML priority after follow-through (dark blue) divided by average rule priority (light blue) yields the relative trust demonstrated in ML after follow-through (mid-blue). Horizontal lines are averages.
@@
Our ML models were designed to be trustworthy, so the outstanding question is whether there is a relationship between the models’ prioritisation of recommendations and the outcomes delivered, i.e. did participants who implemented high-scoring interventions get better outcomes? Was the risk worth it? The Deserving Of Trust index (DOTI)measures whether the trust was worth it, reflecting the constraints faced once the participant acted. There will be a placebo effect at play, which can be considerable, but there is no reason to assume it would not align to demonstrated trust.
Demonstrated trust is shown on an absolute basis on the left of Figure 11 – i.e. the correlation (${r}^{2}$) between each individual’s average ML or rules score and their percentage increase in TST and SWS. We compare baseline to the last seven days of intervention because most participants worked towards target behaviour in stages over four weeks. In this pilot, the strongest relationship was for ML-interventions at follow-through (dark blue): TST ${r}^{2}$ increases from 0.58 at intention to 0.76; SWS from 0.22 to 0.54. An increase is consistent with a trustworthy model because recommendations only work if people act on them. Relative DOTI, the ratio of the absolute ML score and the rules score, is shown on the right. Trust was more deserved in the ML algorithm for TST than for SWS, which may have been influenced by lower SWS model accuracy. At follow-through, our ML model deserved more trust than the rules-based one; had only intention been considered, the SWS rules-based model would have (incorrectly) been seen to deserve more trust.
Figure 11: Absolute DOTI on the left: TST at top and SWS on bottom: correlation (${r}^{2}$) between intention and follow-through scores, and % improvement in last 7 days’ sleep end points relative to baseline. Relative DOTI is on the right (ML metric divided by rules-based).
@@ 5 Discussion
The proposed architecture allows the ML leap of faith to be measured and managed, contributing guidance on how to:
  • Build user intrinsic trust in AI agents, by pre-validating data in an accessible format, and deciding their objectives. This functional compartmentalisation isolates algorithmic reasoning.
@@
  • Build expert intrinsic trust, by providing a rules-based model that they test a priori against their mental model. If this were collectively-developed, it could become a reference agent. Industry or regulatory bodies could maintain the agent as critical infrastructure – e.g. for assurance audits or to train up new experts. Many domains already use rules-based models (e.g. banking), which could become reference standards.
@@
  • Bring the ML leap of faith into sharp focus: a LoF matrix visually and accessibly contrasts the output of the reference agent with the ML model’s a posteriori output, using the same data and objective function. This provides real transparency to system users: they should be inclined to trust where agents agree, and understand a leap of faith is only required where they disagree. Experts can stress test the areas of disagreement, updating the rules-based agent with new insights, thus narrowing the leap of faith. Different ML algorithms will require different leaps of faith, because a priori levels of expert insight differ, as do model accuracies.
@@
  • Assess demonstrated and deserved trust, by considering recommendation actions and outcomes. If trust is deserved, so relative DOTI increases over time; DIRTI and DAFTI should also increase. These simple metrics allow trust comparisons across models, and flag the need for investigation if they trend down. Industry regulators could track companies’ ML-model-trust metrics, especially DOTI, to understand trends.
Limitations include:
@@
  • The matrix and metrics may be perceived as insufficiently technical. Rather, they are accessible by design, so they can be used and understood by experts, businesses, and regulators as well as computer/ data scientists. They need to be applied at scale to reach empirical conclusions about trust, and what drives it, objectively.
  • The proposed metrics do not measure attitudinal trust or model trustworthiness. Instead they measure how these are reflected in the drivers of ML adoption: decisions, actions, and outcomes. If GDPR (or similar) is not followed, trust metrics could be inappropriately sold and/ or used to discriminate based on people’s trust inclination.
@@
  • Participant time for data validation can be substantial, especially if data does not readily align with perception. Our approach may be most suitable for professionals who use agent recommendations often, so their time commitment pays off as they focus on where there is a leap of faith to be overcome. The data-validation approach is not universally accessible: our participants had at least one university degree and understood complex, novel concepts and images. Work is needed to evaluate accessibility for other levels of educational attainment.
  • The time required to develop a reference agent from scratch can be considerable, requiring: substantial programmer subject-matter education; individuals who can translate expert knowledge into rules; and time from experts on each iteration, and to test the resulting agent for different use cases.
  • Prioritisation, unlike direct regression/ classification comparisons, requires calibration if it is to be included in a LoF matrix: different approaches should be tested.
@@ Conclusion
The proposed architecture is a more practical, rigorous way of building trust in an ML agent than trying to explain it to people who have a different (probably unarticulated) mental model. Critical novel architecture features include: user data validation and objective-function specification; an explicit reference standard agent; a LoF matrix; and demonstrated- and deserved-trust metrics.